I have been working on this equation involving 2 radicals for 2 days. Here is what I have so far.
radical(3x+1)-1- radical (x+4) = 0
isolating 1 radical: radical (3x+1)-1 = radical (x+4)
squaring both sides of the equation: (radical(3x+1)-1)^2 = (radical(x+4))^2
(3x+1)+ 2 radical(3x+1)+1 = x+4
2 radical(3x+1) = -2x+2
squaring both sides again: (2 radical(3x+1))^2 = (-2x+2)^2
4(9x+1)=4x^2-4x+4
36x+4= 4x^2-4x+4
factoring:0=4x^2-40x
Where do I go from here to get the answer of 5 as a solution?

Answer by scott8148(6628)

there is a sign error in the first squaring step

(3x+1)(-) 2 radical(3x+1)+1 = x+4

-2 radical(3x+1) = -2x+2 ___ dividing by -2

radical(3x+1) = x - 1 ___ squaring

3x + 1 = x^2 - 2x + 1 ___ 0 = x^2 - 5x ___ 0 = x(x - 5)